Number 1
riple of U.s you that the l hours Ch. 11 Test Review A normal distribution has mean u and distribution. Find the following probabi For 4-6, a normal distribution has a mea selected x-value from the distribution i
Answer #1

P(\mu - 2\sigma< x < \mu + \sigma )

= P((\mu - 2\sigma - \mu)/\sigma< (X - \mu )/\sigma< (\mu + \sigma - \mu)/\sigma)

= P(-2 < Z < 1)

= P(Z < 1) - P(Z < -2)

= 0.8413 - 0.0228

= 0.8185